OpenStudy (anonymous):

The coefficient of x in the expansion of (x+(1/a(x)^2)^7 is 7/3. find the possible values of a

OpenStudy (anonymous):

the power 7 on all expansion or on the next term

OpenStudy (anonymous):

if the expansion is (x+(1/a(x)^2))^7 the expand is

OpenStudy (anonymous):

put coefficient of x=7/3 21/a=7/3 a=9

OpenStudy (anonymous):

only 9? can a also equal -9?

OpenStudy (anonymous):

there exist something wrong

OpenStudy (anonymous):

in this expansion no coefficient to x